One week a business sold 40 scarves. White ones cost $4.95, and printed ones cost $7.95. In all,$282 worth of scarves were sold. How many of each kind were sold?
Answer by checkley77(12844)

w+p=40 or w=40-p
4.95w+7.95p=282
4.95(40-p)+7.95p=282
198-4.95p+7.95p=282
3p=282-198
3p=84
p=84/3
p=28 printed scarves were sold.
40-28=12 white scarves were sold.
Proof:
4.95*12=7.95*28=282
59.4+222.6=282
282=282